1)If P>Q =4i+3j+2k
Find the magnitude of PQ' i.e /PQ/
(2) if A'=3i-j+2k and'B=2i+3j-k, Find
(1) 'A×B'
(2) A'×(2A'+3B')
(3)A'.B'
1
Expert's answer
2021-07-30T11:29:01-0400

Assume

"P=4\hat{i}+3\hat{j}+2\hat{k}"

"Q=2\hat{i}+3\hat{j}+\hat{k}"

"|P|=\sqrt{16+9+4}=\sqrt{30}\\|Q|=\sqrt{4+9+1}=\sqrt{14}"

"P>Q" correct statement

PQ="(4\hat{i}+3\hat{j}+2\hat{k}).(2\hat{i}+3\hat{j}+\hat{k})"

PQ=8+9+2=19

2.

"A=3\hat{i}-\hat{j}+2\hat{k}\\B=2\hat{i}+3\hat{j}-\hat{k}"

"A\times B=\begin{vmatrix}
 \hat{i} & \hat{j}&\hat{k} \\
 3& -1&2\\2&3&-1
\end{vmatrix}"


"A\times B=\hat{i}(1-6)-\hat{j}(-3-4)+\hat{k}(9+2)"

"A\times B=-5\hat{i}+7\hat{j}+11\hat{k}"

(2)


"2A+2B=2\times(3\hat{i}-\hat{j}+2\hat{k})+2\times (2\hat{i}+3\hat{j}-\hat{k})"

"(2A+2B)=10\hat{i}+4\hat{j}+2\hat{k}"

"A\times(2A+2B)=\begin{vmatrix}
 \hat{i} & \hat{j}&\hat{k} \\
 3& -1&2\\10&4&2
\end{vmatrix}"

"A\times(2A+2B)=\hat{i}(-2-6)-\hat{j}(6-20)+\hat{k}(12+10)=-8\hat{i}+14\hat{j}+22\hat{k}"

"A.B=(3\hat{i}-\hat{j}+2\hat{k}).(2\hat{i}+3\hat{j}-\hat{k})=6-6-2=-2"